# Chiral dynamics with giant systems

### v1

@@ Abstract
By fixing $\theta_{Rk_{i}}=\pi-\varphi$, we investigate the statistics of transmitted photons for right-moving inputs, varying both $k_{i}$ and $\varphi$. The results are depicted in Fig. 2. When $U$ is comparable to the decay rate $\gamma$, the output photons can exhibit sub-Poissonian ($g_{R}^{(2)}(0)<1$) or super-Poissonian ($g_{R}^{(2)}(0)>1$) statistics for input momenta $k_{i}$ within a width approximately $\gamma$. In contrast to the single-point coupling scenario Shi et al. (2011), nontrivial statistical behaviors of transmitted photons persist even in the strong dissipation regime where $\gamma\gg U$, as shown in Fig. 2(b). Further tuning of $k_{i}$ and $\varphi$ reveals that the second-order correlation function $g_{R}^{(2)}(0)$ varies in a wide range from $10^{-2}$ to $10^{4}$ (Fig. 2(c)). This nontrivial statistical behavior arises from photon interference between the two coupling points, resulting in an effective decay rate $-\text{Im}[\Sigma(k)]=2\gamma[1+\cos\varphi\cos\theta_{Rk}]$ for the $k$-th mode, which can be significantly smaller than $U$. These modes can be excited in the two-photon scattering process by matching the input momentum $k_{i}$ with the real part of the its self-energy, $\text{Re}[\Sigma(k)]=2\gamma\cos\varphi\sin\theta_{k}$. Fig. 2(d) illustrates this phenomenon with the background fluorescence Shen and Fan (2007); Shi and Sun (2009) $\Gamma_{p_{2}p_{1},k_{i}k_{i}}$ at the point indicated by the rectangle in Fig. 2(c), where $k_{i}=0.09\gamma$, $\varphi=0.015\pi$, and $g_{R}^{(2)}(0)=0.017$, corresponding to $\Sigma(k_{i})=(0.094-i0.0044)\gamma$. Fig. 2(d) shows highly populated modes centered at $k_{i}+n\frac{2\pi}{d}$ with a narrow width of --Im$[\Sigma(k)]/\left(\gamma L\right)$, underscoring the nontrivial statistical behaviors of transmitted photons.
Figure 2: Properties of the output photons in the weak-driving limit. (a) Second-order correlation function $g_{R}^{(2)}(0)$ of the transmitted light with $\phi=\pi/15$ and $U=\gamma$. (b) Same as (a), but for weak nonlinearity $U=0.1\gamma$. (c) $g_{R}^{(2)}(0)$ as a function of $k_{i}$ and $\phi$ in the strong dissipation limit with $U=0.01\gamma$. (d) Background fluorescence $\left|\Gamma_{p_{2}p_{1},k_{i}k_{i}}\right|$ with parameters from the rectangle in (c). Here, in $\Gamma_{p_{2}p_{1},k_{i}k_{i}}$, $p_{2}=2k_{i}-p_{1}$.
@@
When $\gamma d\ll 1$, the two-photon process simplifies to be Markovian (dashed-dotted line in Figs. 2(a) and (b)), and the system’s evolution can be approximatly described by a master equation and the input-output formalism Gardiner and Zoller (2004); Walls and Milburn (2008). Generally, this Markovian approximation breaks down for large distances $d$, as evidenced by the deviation in the second-order correlator $g_{R}^{(2)}(0)$ between $\gamma d=0.01$ and $\gamma d=1,100$ (for fixed $\theta_{Rk_{i}}$, Markovian dynamics are independent of $d$). However, when $\phi=\pm\pi/2$, the self-energy $\Sigma(k)$ becomes independent of $k$, and the evolution of the giant cavity’s density matrix $\rho$ can be exactly described by the master equation Shi et al. (2015) regardless of the distance $d$ :
@@
where $\rho$ is the density matrix of the cavity and $H_{S}=\frac{U}{2}b^{{\dagger}2}b^{2}+H_{\lambda D}$. In this scenario, the decay of the cavity due to the interference between right- and left-moving photons propagating between the two coupling points is completely canceled out. This cancellation can also be understood by defining two modes
@@
where the unitary transformation $U_{k}=\frac{1}{\tilde{V}_{k}}\left(\begin{array}[]{cc}V_{L-k}&V_{Rk}\\ V_{Rk}^{\ast}&-V_{L-k}^{\ast}\end{array}\right)$, with $\tilde{V}_{k}=2V_{0}\sqrt{1+\cos\phi\cos\left(k_{0}+k\right)L}$. These two modes are degenerated and decoupled, and the cavity field only couples to $A_{+k}$ mode with the interaction in Eq. (2) rewritten as
@@
When $\phi=\pm\pi/2$, $\tilde{V}_{k}$ is independent of $k$, leading to a constant decay rate $2\gamma$ contributed from the incoherent superposition of cavity photons decaying via the two coupling points. Note that this $k$-independence in $\tilde{V}_{k}$ with $\phi=\pm\pi/2$ is general and does not depend on the specific spectrum $\varepsilon_{\lambda k}$. However, the Markovian description is rigorously valid only for a linear spectrum. Furthermore, the properties of photons propagating in the waveguide can be acquired with the input-output relation
@@
$a_{out}^{\lambda}\left(t\right)=a_{in}^{\lambda}\left(t\right)-i\sqrt{\gamma}\left[e^{-i\phi}b\left(t\right)+e^{-i\theta_{\lambda 0}}b\left(t+\sigma_{\lambda}d\right)\right],$ (13)
@@
where the input field $a_{in}^{\lambda}\left(t\right)=\frac{1}{\sqrt{2\pi}}\int a_{\lambda k}\left(0\right)e^{-i\sigma_{\lambda}kt}\mathrm{d}k$. Note that unlike the giant system itself, the output field $a_{out}^{\lambda}\left(t\right)$ is determined by the cavity field at two distinct times $t$ and $t\pm d$ even when $\phi=\pm\pi/2$. This indicates that the non-Markovian characteristic is preserved in the properties of the output photons.
@@
The master equation (10), along with the input-output formalism (13), enables a comprehensive exploration of the giant system’s dynamics and the characteristics of the output photons under arbitrarily strong driving fields. When $\phi=(k_{0}+k_{i})d=\pi/2$, the giant system interacts exclusively with the right-moving driving field, leading to the emergence of a DPT in the strong-driving regime Drummond and Walls (1980); Drummond and Gardiner (1980); Bartolo et al. (2016). As depicted in Fig. 3(a), increasing the pump strength $\Omega_{0}$ results in a sudden increase in the cavity photon number $\langle b^{{\dagger}}b\rangle_{S}$ around $\Omega_{0}\approx 5\gamma$, corresponding to a peak in the second-order correlator $g^{(2)}(0)\equiv\frac{\langle b^{{\dagger}2}b^{2}\rangle_{S}}{\langle b^{{\dagger}}b\rangle_{S}^{2}}$, where $\langle...\rangle_{S}$ denotes the steady-state average. Here, we consider Kerr nonlinearity $U=\gamma$ and input momentum $k_{i}=10\gamma$. Using methods such as the complex P-representation formalism Walls and Milburn (2008); Drummond and Walls (1980); Drummond and Gardiner (1980); Bartolo et al. (2016) or the quantum absorber approach Stannigel et al. (2012); Roberts and Clerk (2020), one can analytically derive the steady-state solution of the master equation (10), revealing a first-order DPT in the thermodynamic limit where the Liouvillian gap (the real part of the second-largest eigenvalue of the Liouvillian) vanishes at the critical point Minganti et al. (2018). Remarkably, this DPT-induced sudden change manifests not only in the cavity field but also in the reflected field, highlighting the non-Markovian effects induced by the giant system. At $T\rightarrow\infty$, the density $\rho_{L}\equiv\frac{1}{2\pi}\int\langle a_{Lk_{2}}^{{\dagger}}(T)a_{Lk_{1}}(T)\rangle e^{i(k_{1}-k_{2})X}\mathrm{d}k_{1}\mathrm{d}k_{2}$ of left-moving output photons at position $X<0$ can be computed using the input-output relation (13) as
$\rho_{L}$ $=$ $\left\langle a_{out}^{L{\dagger}}\left(T+x\right)a_{out}^{L}\left(T+x\right)\right\rangle$ (14)
@@
$=$ $2\gamma\left(\left\langle b^{{\dagger}}b\right\rangle_{s}+\text{Im}\left[e^{-ik_{0}d}\left\langle b^{{\dagger}}\left(d\right)b\left(0\right)\right\rangle_{S}\right]\right).$
The reflected photon density $\rho_{L}$ depends on the distance $d$ through the nonlocal correlation $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{S}$. To examine the impact of this two-point correlation, we set $k_{0}d=\pi/2$ and $d=2n\pi/k_{i}$ to satisfy the chirality condition in Eq. (3), and investigate the nonreciprocal dissipative phase transition (DPT) via $\rho_{L}$ for various $d$. When $d=0$, no photons are reflected, while a peak emerges around the critical point as $d$ increases, as illustrated in Fig. 3(b). This peak occurs near the critical point where the Liouvillian exhibits a minimal gap (note that this gap only vanishes in the limit of infinite photon numbers), leading to the maximum correlation $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{S}-\left|\left\langle b\right\rangle_{S}\right|^{2}$ (Fig. 3(b)). As $d$ further increases, the peak in $\rho_{L}$ approaches the critical point, and the magnitude of the correlator $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{S}-\left|\left\langle b\right\rangle_{S}\right|^{2}$ diminishes because only the eigenstate with the smallest gap contributes significantly. Specifically, when $d$ exceeds the Liouvillian gap significantly, such as $d=2000\pi/k_{i}$, represented by the dotted lines in Fig. 3(b), no correlations are present and $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{S}$ approaches $\left|\left\langle b\right\rangle_{S}\right|^{2}$. In this scenario, the photon density $\rho_{L}$ reaches its maximum since $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{S}$ deviates most significantly from $\left\langle b^{{\dagger}}b\right\rangle_{S}$. Therefore, this system provides a platform to investigate the nonlocal effects in the context of DPT.

### v2

@@ II.1 Weak-driving limit
Further tuning of $k_{i}$ and $\varphi$ reveals that the second-order correlation function $g_{\mathrm{R}}^{(2)}(0)$ varies in a wide range from $10^{-2}$ to $10^{4}$ (Fig. 2(c)). This nontrivial statistical behavior arises from photon interference between the two coupling points, resulting in an effective decay rate $\gamma_{k}\equiv$-Im[$\left[\Sigma(k)\right]=2\gamma\left(1+\cos\phi\cos\theta_{\mathrm{R}k}\right)$ for the $k$-th mode, which can be significantly smaller than $U$. These modes can be efficiently excited in the two-photon scattering process by matching its momentum $k$ with the real part of the self-energy, $\text{Re}\left[\Sigma(k)\right]=2\gamma\cos\phi\sin\theta_{\mathrm{R}k}$ or $2k_{i}$-Re[$\left[\Sigma(2k_{i}-k)\right]$. Fig. 2(d) illustrates this phenomenon with the background fluorescence Shen and Fan (2007); Shi and Sun (2009) $\Gamma_{p_{2}p_{1},k_{i}k_{i}}$ at the point indicated by the rectangle in Fig. 2(c), where $k_{i}=0.09\gamma$, $\phi=0.015\pi$, and $g_{\mathrm{R}}^{(2)}(0)=0.017$, corresponding to $\Sigma(k_{i})=(0.094-i0.0044)\gamma$. The blue solid curve in Fig. 2(d) shows highly populated modes centered around $0.094+2n\pi/d$ with a narrow width of $\gamma_{k}/\left(\gamma L\right)$, underscoring the nontrivial statistical behaviors of transmitted photons. In contrast, when $d$ decreases, the number of modes satisfying k=k=Re[$\left[\Sigma(k)\right]$ decrease, leading to less resonant peaks in the background fluorescence, together with larger width $\gamma_{k}$. Specifically, as shown in the red dashed curve, in the Makovian regime with $\gamma d=0.01$, the equation k=k=Re[$\left[\Sigma(k)\right]$ has only one resonant peak, resulting in different statistics of output photons.
We note that throughout the manuscript, the frequency of the cavity is set equal to the central frequency $k_{0}$ of the waveguide. In a gernal case, a detuning $\Delta_{c}$ between the cavity frequency and $k_{0}$ can be introduced. With this detuning, all occurrences of $k$ in the expressions—except in the phase factor $\theta_{\lambda k}$—should be replaced by $k-\Delta_{\mathrm{c}}$. By redefining $\tilde{k}=k-\Delta_{\mathrm{c}}$, the phase $\theta_{\mathrm{R}k}=\left(k_{0}+\tilde{k}_{i}+\Delta_{c}\right)d$, and the chirality condition in Eq. (3) is modified to
$\phi=\left(2n+1\right)\mathrm{\pi}-\left(k_{0}+\tilde{k}_{i}+\Delta_{c}\right)d,$ (14)
Since $\theta_{\mathrm{R}k}=\left(2n+1\right)\pi-\phi$, this redefinition allows all other expressions to retain their original form with $k$ replaced directly by $\tilde{k}$. Thus, when specifying $\tilde{k}_{i}$ as the input momentum, the results remain unchanged. For given values of $\phi$, $k_{0}$, and $d$, the chirality condition can be satisfied for different $\tilde{k}_{i}$ by tuning $\Delta_{\mathrm{c}}$.
II.2 Markovianity at $\phi=\pm\mathrm{\pi}/2$
@@
When $\gamma d\ll 1$, the two-photon dynamics simplify to be Markovian (dashed-dotted line in Figs. 2(a) and (b)), and the system’s evolution can be approximatly described by a master equation and the input-output formalism Gardiner and Zoller (2004); Walls and Milburn (2008). Generally, this Markovian approximation breaks down for large distances $d$, as evidenced by the deviation in the second-order correlator $g_{\mathrm{R}}^{(2)}(0)$ between $\gamma d=0.01$ and $\gamma d=1,100$ shown in Figs. 2(a) and (b) (for fixed $\theta_{\mathrm{R}k_{i}}$, Markovian dynamics are independent of $d$). However, when $\phi=\pm\mathrm{\pi}/2$, the self-energy $\Sigma(k)$ becomes independent of $k$, and the evolution of the giant cavity’s density matrix $\rho$ can be exactly described by the master equation Shi et al. (2015) regardless of the distance $d$ and photon number:
@@
where $\rho$ is the density matrix of the cavity and $H_{S}=\frac{U}{2}b^{{\dagger}2}b^{2}+H_{\lambda D}$. In this scenario, the decay of the cavity due to the interference between right- and left-moving photons propagating between the two coupling points is completely canceled out.
This cancellation can also be understood by defining two modes
@@
where the unitary transformation $U_{k}=\frac{1}{\tilde{V}_{k}}\left(\begin{array}[]{cc}V_{\mathrm{L}-k}&V_{\mathrm{R}k}\\ V_{\mathrm{R}k}^{\ast}&-V_{\mathrm{L}-k}^{\ast}\end{array}\right)$, with $\tilde{V}_{k}=2V_{0}\sqrt{1+\cos\phi\cos\left(k_{0}+k\right)L}$. These two modes are degenerated and decoupled, and the cavity field only couples to $A_{+k}$ mode with the interaction in Eq. (2) rewritten as
@@
When $\phi=\pm\mathrm{\pi}/2$, $\tilde{V}_{k}$ is independent of $k$, leading to a constant decay rate $2\gamma$ contributed from the incoherent superposition of cavity photons decaying via the two coupling points. Note that this $k$-independence in $\tilde{V}_{k}$ with $\phi=\pm\pi/2$ is general and does not depend on the specific spectrum $\varepsilon_{\lambda k}$. However, the Markovian description is rigorously valid only for a linear spectrum. Furthermore, the properties of photons propagating in the waveguide can be acquired with the time-nonlocal input-output relation
@@
$a_{\mathrm{out}}^{\lambda}\left(t\right)=a_{\mathrm{in}}^{\lambda}\left(t\right)-i\sqrt{\gamma}\left[e^{-i\phi}b\left(t\right)+e^{-i\theta_{\lambda 0}}b\left(t+\sigma_{\lambda}d\right)\right],$ (18)
@@
where the input field $a_{\mathrm{in}}^{\lambda}\left(t\right)=\frac{1}{\sqrt{2\pi}}\int a_{\lambda k}\left(0\right)e^{-i\sigma_{\lambda}kt}\mathrm{d}k$. Note that unlike the giant system itself, the output field $a_{\mathrm{out}}^{\lambda}\left(t\right)$ is determined by the cavity field at two distinct times $t$ and $t\pm d$ even when $\phi=\pm\mathrm{\pi}/2$. This indicates that the non-Markovian characteristic is preserved in the properties of the output photons.
II.3 Non-Makovian dissipative phase transition
@@
The master equation (15), along with the input-output formalism (18), enables a comprehensive exploration of the giant system’s dynamics and the characteristics of the output photons under arbitrarily strong driving fields. When $\phi=(k_{0}+k_{i})d=\mathrm{\pi}/2$, the giant system interacts exclusively with the right-moving driving field, leading to the emergence of a DPT in the strong-driving regime Drummond and Walls (1980); Drummond and Gardiner (1980); Bartolo et al. (2016). As depicted in Fig. 3(a), increasing the pump strength $\Omega_{0}$ results in a sudden increase in the cavity photon number $\langle b^{{\dagger}}b\rangle_{\mathrm{S}}$ around $\Omega_{0}\approx 5\gamma$, corresponding to a peak in the second-order correlator $g^{(2)}(0)\equiv\frac{\langle b^{{\dagger}2}b^{2}\rangle_{\mathrm{S}}}{\langle b^{{\dagger}}b\rangle_{\mathrm{S}}^{2}}$, where $\langle...\rangle_{\mathrm{S}}$ denotes the steady-state average. Here, we consider the Kerr nonlinearity $U=\gamma$ and incident momentum $k_{i}=10\gamma$. Using methods such as the complex P-representation formalism Walls and Milburn (2008); Drummond and Walls (1980); Drummond and Gardiner (1980); Bartolo et al. (2016) or the quantum absorber approach Stannigel et al. (2012); Roberts and Clerk (2020), one can analytically derive the steady-state solution of the master equation (15), revealing a first-order DPT in the thermodynamic limit where the Liouvillian gap (the real part of the second-largest eigenvalue of the Liouvillian) vanishes at the critical point Minganti et al. (2018).
This DPT-induced sudden change manifests not only in the cavity field but also in the reflected field, highlighting the non-Markovian effects induced by the giant system. At $T\rightarrow\infty$, the density $\rho_{L}\equiv\frac{1}{2\mathrm{\pi}}\int\langle a_{Lk_{2}}^{{\dagger}}(T)a_{Lk_{1}}(T)\rangle e^{i(k_{1}-k_{2})X}\mathrm{d}k_{1}\mathrm{d}k_{2}$ of left-moving output photons at position $X<0$ can be computed using the input-output relation (18) as
$\rho_{L}$ $=$ $\left\langle a_{\mathrm{out}}^{\mathrm{L}{\dagger}}\left(T+x\right)a_{\mathrm{out}}^{\mathrm{L}}\left(T+x\right)\right\rangle$ (19)
@@
$=$ $2\gamma\left(\left\langle b^{{\dagger}}b\right\rangle_{\mathrm{S}}-\text{Im}\left[e^{ik_{0}d}\left\langle b^{{\dagger}}\left(d\right)b\left(0\right)\right\rangle_{\mathrm{S}}\right]\right).$
The DPT observed in the reflected field can differ significantly from that of the cavity field. For instance, under the conditions $\phi=\pm\pi/2$ and Eq. (3), the phase transition in the cavity field is independent of $d$. In contrast, the critical behavior in the reflected field is highly sensitive to $d$ due to the nonlocal correlation $\left\langle b^{{\dagger}}(d)b(0)\right\rangle_{\mathrm{S}}$. When $d$ approaches zero, no reflected photons are observed, and thus no phase transition occurs in the reflected field. As $d$ increases, the DPT in the reflected field emerges, revealing the non-Markovian nature of the critical behavior in the output field.